Charge is distributed on the surface of a sphere with a surface charge density σ=σ0 cos θ in spherical coordinates. This distribution is


(a) Spherically Symmetric


(b) cylindrically symmetric


(c) Planar Symmetric


(d) none of these


(Note:- This question have one or more than one correct choice(s) out of the four given choices. Any number of options may be correct.)


1
Expert's answer
2020-09-22T15:35:47-0400

We can represent spherical coordinates in cartesian as following



(picture from Wikipedia). We can see here that "cos \theta" gives us projection on z-axis. So all values of "\u03c3_0 cos \u03b8" are equal for the same value of theta. Cosine is an even function, "cos\theta = cos(\pi-\theta)"

So, values are the same for +z and -z. This gives us planar symmetry relative to the plane "\theta =0" or X0Y in cartesian coordinates.

Answer: (c) Planar Symmetric